$SPAD/src/input f02asf.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1994.

———-

— * —

\getchunk{license}
)cl all
showArrayValues true
showScalarValues true
ia:=8
b:Matrix SF:=
 [[23 ,32 ,33 ,31 ]]
n:=4
a:Matrix SF:=
 [[5,7,6,5],_
  [7,10,8,7],_
  [6,8,10,9],_
  [5,7,9,10],_
  [0,0,0,0],_
  [0,0,0,0],_
  [0,0,0,0],_
  [0,0,0,0]]
result:=f04asf(ia,b,n,a,-1)

———-

References^{[1] nothing}

4